# D2-VPR: A Parameter-efficient Visual-foundation-model-based Visual Place Recognition Method via Knowledge Distillation and Deformable Aggregation

### v3

@@ Abstract
Visual Place Recognition (VPR) aims to determine the geographic location of a query image by retrieving its most visually similar counterpart from a geo-tagged reference database. Recently, the emergence of the powerful visual foundation model, DINOv2, trained in a self-supervised manner on massive datasets, has significantly improved VPR performance. This improvement stems from DINOv2’s exceptional feature generalization capabilities but is often accompanied by increased model complexity and computational overhead that impede deployment on resource-constrained devices. To address this challenge, we propose $D^{2}$-VPR, a DDistillation- and DDeformable-based framework that retains the strong feature extraction capabilities of visual foundation models while significantly reducing model parameters and achieving a more favorable performance-efficiency trade-off. Specifically, first, we employ a two-stage training strategy that integrates knowledge distillation and fine-tuning. Additionally, we introduce a Distillation Recovery Module (DRM) to better align the feature spaces between the teacher and student models, thereby minimizing knowledge transfer losses to the greatest extent possible. Second, we design a Top-Down-attention-based Deformable Aggregator (TDDA) that leverages global semantic features to dynamically and adaptively adjust the Regions of Interest (ROI) used for aggregation, thereby improving adaptability to irregular structures. Extensive experiments demonstrate that our method achieves competitive performance compared to state-of-the-art approaches. Meanwhile, it reduces the parameter count by approximately 64.2% and MACs by about 62.6% (compared to CricaVPR).
@@ Introduction
Figure 1: The comparison of average R@5 against multiply-accumulate operations (MACs) and parameter count on Pitts30k, MSLS-val, and SPED (with image size of 224×224) demonstrates that our model achieves competitive performance despite significantly reduced MACs and parameter count, striking an effective trade-off.
@@
To summarize, our work makes the following contributions: 1) We have designed a two-stage training strategy that combines knowledge distillation and fine-tuning to train a parameter-efficient VPR model based on visual foundation models. Additionally, we propose the Distillation Recovery Module (DRM) to minimize knowledge loss during the knowledge distillation process to the greatest extent possible. 2) We design a Top-Down-attention-based Deformable Aggregator (TDDA), which controls the deformation of local ROIs through global semantic information and demonstrates better adaptability to irregular structures and regions compared with existing aggregators centered on spatial pooling. 3) Extensive experiments show that our $D^{2}$-VPR can deliver competitive performance compared with SOTA methods on popular benchmarks while achieving significant reductions in parameter count, MACs, and storage consumption.
@@
Method Dim. Dinov2 Param. MACs
(M) (G)
CricaVPR $_{{\text{CVPR' 2024}}}$ 4096 Base 106.76 24.41
SALAD $_{{\text{CVPR' 2024}}}$ 8448 Base 87.99 22.19
BoQ $_{{\text{CVPR' 2024}}}$ 12288 Base 95.21 22.64
SuperVLAD $_{{\text{NIPS' 2024}}}$ 3072 Base 97.61 21.99
Sela-global $_{{\text{ICLR' 2024}}}$ 1024 Large 357.43 94.14
FoL-global $_{{\text{AAAI' 2025}}}$ 8448 Large 308.83 78.92
EDTformer $_{{\text{TCSVT' 2025}}}$ 4096 Base 96.96 22.15
$D^{2}$-VPR $_{{\text{no encoder}}}$ 4096 Small 27.21 9.05
$D^{2}$-VPR $_{{\text{encoder}}}$ 4096 Small 38.24 9.14
@@
Comparison with Larger-scale SOTA Baselines. We also comprehensively compare our proposed method with SOTA one-stage VPR methods in Tables 4 and 5, including CricaVPR (crica), SALAD (salad), BoQ (ali2024boq), SuperVLAD (lu2024supervlad), SelaVPR (sela), FoL (wang2025focus), and EDTformer (EDTformer). Note that we have unified the evaluation image resolution to 224×224 rather than a higher resolution, which aligns with our goal of deploying the model on resource-constrained devices. When using the cross-image encoder, the results demonstrate competitive performance, particularly on Pitts30k, MSLS-val, Pitts250k, and AmsterTime. Figure 4 provides visualizations comparing the retrieval performance of our method with other baselines. However, due to the inherent limitations of the cross-image encoder, we also report comparisons without it. In this setting, our method achieves performance comparable to SALAD on Pitts30k, MSLS-val, and Pitts250k, and performs better than FoL-global on Nordland, indicating that our approach still maintains a reasonable level of competitiveness, despite requiring substantially fewer parameters and MACs than the competing methods.

### v4

@@ Abstract
Visual Place Recognition (VPR) aims to determine the geographic location of a query image by retrieving its most visually similar counterpart from a geo-tagged reference database. Recently, the emergence of the powerful visual foundation model, DINOv2, trained in a self-supervised manner on massive datasets, has significantly improved VPR performance. This improvement stems from DINOv2’s exceptional feature generalization capabilities but is often accompanied by increased model complexity and computational overhead that impede deployment on resource-constrained devices. To address this challenge, we propose $D^{2}$-VPR, a DDistillation- and DDeformable-based framework that retains the strong feature extraction capabilities of visual foundation models while significantly reducing model parameters and achieving a more favorable performance-efficiency trade-off. Specifically, first, we employ a two-stage training strategy that integrates knowledge distillation and fine-tuning. Additionally, we introduce a Distillation Recovery Module (DRM) to better align the feature spaces between the teacher and student models, thereby minimizing knowledge transfer losses to the greatest extent possible. Second, we design a Top-Down-attention-based Deformable Aggregator (TDDA) that leverages global semantic features to dynamically and adaptively adjust the Regions of Interest (ROI) used for aggregation, thereby improving adaptability to irregular structures. Extensive experiments demonstrate that our method achieves competitive performance compared to state-of-the-art approaches. Meanwhile, it reduces the parameter count by approximately 64.2% (compared to CricaVPR).
@@ Introduction
Figure 1: The comparison of average R@5 against parameter count on Pitts30k, MSLS-val, and SPED (with image size of 224×224) demonstrates that our model achieves competitive performance despite significantly reduced parameter count, striking an effective trade-off.
@@
To summarize, our work makes the following contributions: 1) We have designed a two-stage training strategy that combines knowledge distillation and fine-tuning to train a parameter-efficient VPR model based on visual foundation models. Additionally, we propose the Distillation Recovery Module (DRM) to minimize knowledge loss during the knowledge distillation process to the greatest extent possible. 2) We design a Top-Down-attention-based Deformable Aggregator (TDDA), which controls the deformation of local ROIs through global semantic information and demonstrates better adaptability to irregular structures and regions compared with existing aggregators centered on spatial pooling. 3) Extensive experiments show that our $D^{2}$-VPR can deliver competitive performance compared with SOTA methods on popular benchmarks while achieving significant reductions in parameter count and storage consumption.
@@
Method Dim. Dinov2 Image Param.
Size (M)
CricaVPR $_{{\text{CVPR' 2024}}}$ 4096 Base 224×224 106.76
SALAD $_{{\text{CVPR' 2024}}}$ 8448 Base 224×224 87.99
BoQ $_{{\text{CVPR' 2024}}}$ 12288 Base 224×224 95.21
SuperVLAD $_{{\text{NIPS' 2024}}}$ 3072 Base 224×224 97.61
Sela-global $_{{\text{ICLR' 2024}}}$ 1024 Large 224×224 357.43
FoL-global $_{{\text{AAAI' 2025}}}$ 8448 Large 224×224 308.83
EDTformer $_{{\text{TCSVT' 2025}}}$ 4096 Base 224×224 96.96
$D^{2}$-VPR $_{{\text{no encoder}}}$ 4096 Small 224×224 27.21
$D^{2}$-VPR $_{{\text{encoder}}}$ 4096 Small 224×224 38.24
@@
Comparison with Larger-scale SOTA Baselines. We also comprehensively compare our proposed method with SOTA one-stage VPR methods in Tables 4 and 5, including CricaVPR (crica), SALAD (salad), BoQ (ali2024boq), SuperVLAD (lu2024supervlad), SelaVPR (sela), FoL (wang2025focus), and EDTformer (EDTformer). Note that we have unified the evaluation image resolution to 224×224 rather than a higher resolution, which aligns with our goal of deploying the model on resource-constrained devices. When using the cross-image encoder, the results demonstrate competitive performance, particularly on Pitts30k, MSLS-val, Pitts250k, and AmsterTime. Figure 4 provides visualizations comparing the retrieval performance of our method with other baselines. However, due to the inherent limitations of the cross-image encoder, we also report comparisons without it. In this setting, our method achieves performance comparable to SALAD on Pitts30k, MSLS-val, and Pitts250k, and performs better than FoL-global on Nordland, indicating that our approach still maintains a reasonable level of competitiveness, despite requiring substantially fewer parameters than the competing methods.